OpenStudy (anonymous):

the position of a particle moving along a horizontal line is given by x(t)=4 (t-4)^3. What is the maximum speed of the particle for 0

OpenStudy (anonymous):

find the max of the derivative

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

would the derivative be x'(t)=12 (t-4)^2

OpenStudy (anonymous):

that's right..

OpenStudy (anonymous):

v(t)=x'(t)=12 (t-4)^2

OpenStudy (anonymous):

this is a parabola that faces up so it doesn not have a maximum however, your interval is \(0\leq t\leq 10\) so you only have to check at the endpoints for a max

OpenStudy (anonymous):

t=10 v=432 m/s